OpenStudy (anonymous):

If G(x) = 3x + 1, then G-1(1) is A. 4 B. 1/3 C. 0

OpenStudy (anonymous):

C

OpenStudy (matt101):

G-1(x)=(x-1)/3 G-1(1)=0